# Interpreting Global Perturbation Robustness of Image Models using Axiomatic Spectral Importance Decomposition

### v2

@@ 3.2 Spectral coalition filtering
To implement the presences and absences of spectral signals, we define a mask map $\mathbbm{T}(\widetilde{\mathcal{I}}):\widetilde{\mathcal{I}}\mapsto\{0,1\}^{M\times N}$ on 2D spectrum, where $\widetilde{\mathcal{I}}$ is a spectral coalition and $M\times N$ denotes 2D image dimensions. The mask map is point-wisely defined as:
$\mathbbm{T}(\widetilde{\mathcal{I}})(m,n)=\begin{cases}1,&\text{if the frequency point}\leavevmode\nobreak\ $(m,n)$\leavevmode\nobreak\ \text{is present in coalition}\leavevmode\nobreak\ \widetilde{\mathcal{I}},\\ 0,&\text{otherwise}\end{cases}$ (4)
@@
We apply the transfer function $\mathbbm{T}(\widetilde{\mathcal{I}})$ on the spectra of images with element-wise product (i.e. Hadamard product (Horn, 1990; Horadam, 2012)). Let $\mathscr{F}$ be the Discrete Fourier transform (DFT) operator and $\mathscr{F}^{-1}$ be the inverse DFT (IDFT) operator (Tan & Jiang, 2018). Readers can further refer to Appendix B.1.
@@
$x\bowtie\widetilde{\mathcal{I}}:=\mathscr{F}^{-1}\left[\underbrace{\mathscr{F}(x)\odot\mathbbm{T}(\widetilde{\mathcal{I}})}_{\mathrm{Spectral\leavevmode\nobreak\ presence}}+\underbrace{{\bf\it b}\odot(\mathbbm{1}-\mathbbm{T}(\widetilde{\mathcal{I}}))}_{\mathrm{Spectral\leavevmode\nobreak\ absence}}\right]$ (5)
where $\odot$’ denotes Hadamard product (i.e. element-wise product), $\mathbbm{1}\in\mathbb{R}^{M\times N}$ denotes an all-ones matrix and ${\bf\it b}\in\mathbb{C}^{M\times N}$ represents the assignments of the absences of spectral players. In the context of attribution analysis, ${\bf\it b}$ is often referred as the baseline. In our implementation, we empirically set ${\bf\it b}={\bf\it 0}$.
@@ 3.3 Characteristic function design
$v(\widetilde{\mathcal{I}})$ $:=\mathop{\mathbb{E}}\limits_{x\thicksim\mathcal{X}}\sum_{y\in\mathcal{Y}}\left\{\underbrace{P(y|x)\cdot\log Q(y|x\bowtie\widetilde{\mathcal{I}})}_{\mathrm{Feature\leavevmode\nobreak\ presence}}-\underbrace{P(y|x)\cdot\log Q(y|x\bowtie\emptyset)}_{\mathrm{Absence\leavevmode\nobreak\ baseline}}\right\}$
@@
$\mathbb{I}(\mathcal{X}\bowtie\widetilde{\mathcal{I}},\mathcal{Y})\equiv\mathop{\mathbb{missing}}{E}\limits_{x\in\mathcal{X}\bowtie\widetilde{\mathcal{I}}}D_{KL}[P(y|x)||Q(y||x)]+H(\mathcal{Y})+v(\widetilde{\mathcal{I}})+C$ (9)
@@ 3.4 Spectral robustness score (SRS)
The second row in Figure 7 shows the SIDs from various models with randomized weights. We randomize the model weights with Kaiming initialization (He et al., 2015). The measured SIDs exhibit spectral uniformity. This suggests: Un-trained models do not have spectral preferences. We refer to ‘the models with randomized parameters’ as random decisions. Therefore, we assume that the SIDs of random decisions are uniform: $\frac{\mathbbm{1}}{M}$.
@@
|$\absolutevalue{{{\bf\it\beta}}^{T}\bar{\Psi}(v)-{{\bf\it\beta}}^{T}\frac{\mathbbm{1}}{M}}$ | (10)
where ${{\bf\it\beta}}^{T}\frac{\mathbbm{1}}{M}$ is served as a random decision baseline. Let $S(v):v\mapsto[0,1]$ be the normalized result in Equation 10. The $S(v)$ is given by:
$S(v):=\frac{\absolutevalue{{{\bf\it\beta}}^{T}\bar{\Psi}(v)-{{\bf\it\beta}}^{T}\frac{\mathbbm{1}}{M}}}{\sup\limits_{\bar{\Psi}}\absolutevalue{{{\bf\it\beta}}^{T}\bar{\Psi}(v)-{{\bf\it\beta}}^{T}\frac{\mathbbm{1}}{M}}}=\absolutevalue{\frac{{{{\bf\it\bar{\beta}}}}^{T}\bar{\Psi}(v)-\eta}{1-\eta}}$ | (11)
@@ 4 Experiments
(a) White noise ($\sigma=0.1$)
(b) White noise ($\sigma=0.2$)
(c) Gaussian-blur ($3\times 3$)
(d) Gaussian-blur ($7\times 7$)
@@ 4.1 Correlation to robustness metrics
$\Delta\mathcal{P}:=\mathop{\mathbb{missing}}{E}\limits_{x,y\thicksim\langle\mathcal{X},\mathcal{Y}\rangle}|Q(y|x)-Q(y|x^{*})|.$ (12)
@@ Appendix B Fairness division axioms
Symmetry axiom: Let ~$\widetilde{\mathcal{I}}\in 2^{\mathcal{I}}$ be some spectral player coalition. For $\forall\leavevmode\nobreak\ \mathcal{I}_{i},\mathcal{I}_{j}\in\mathcal{I}\land\mathcal{I}_{i},\mathcal{I}_{j}\notin\widetilde{\mathcal{I}}$, the statement $v(\widetilde{\mathcal{I}}\cup\{\mathcal{I}_{i}\})=v(\widetilde{\mathcal{I}}\cup\{\mathcal{I}_{j}\})$ implies $\psi_{i}(\mathcal{I},v)=\psi_{j}(\mathcal{I},v)$. This axiom restates the statement ‘equal treatment of equals’ principle mathematically. This axiom states that the ‘names’ of players should have no effect on the ‘treatments’ by the characteristic function in coalition games (Roth, 1988).
@@
Dummy player axiom: A dummy player (null player) $\mathcal{I}_{*}$ is the player who has no contribution such that: $\psi_{*}(\mathcal{I},v)=0$ and $v(\widetilde{\mathcal{I}}\cup\{\mathcal{I}_{*}\})\equiv v(\widetilde{\mathcal{I}})$ for $\forall\leavevmode\nobreak\ \mathcal{I}_{*}\notin\widetilde{\mathcal{I}}\land\mathcal{I}_{*}\subseteq\mathcal{I}$.
@@ B.1 Spectral signal-to-noise ratio (SNR)
$\mathscr{F}(x)(u,v)=\sum_{j=0}^{N-1}\sum_{i=0}^{M-1}x(i,j)e^{-\mathbbm{i}2\pi(\frac{u}{M}i+\frac{v}{N}j)}.$ (13)
@@ B.2 Absence assignment scheme
(a) Zeroing
(b) Complex Gaussian
(c) Replacement
(d) Zeroing
(e) Complex Gaussian
(f) Replacement
(g) resnet18 w/ Zeroing
(h) resnet18 w/ Complex Gaussian
(i) resnet18 w/ Replacement
(j) efficientnet_v2_s w/ Zeroing
(k) efficientnet_v2_s w/ Complex Gaussian
(l) efficientnet_v2_s w/ Replacement
@@ B.3 Proof for Spectral Coalition Information Identity Theorem
$=\mathop{\mathbb{missing}}{E}\limits_{x\in\mathcal{X}\bowtie\widetilde{\mathcal{I}}}\underbrace{KL(P(y|x)||Q(y||x))}_{\mathrm{point-wise}}+H(\mathcal{Y})+\mathop{\mathbb{missing}}{E}\limits_{x\in\mathcal{X}\bowtie\widetilde{\mathcal{I}}}\sum_{y\in\mathcal{Y}}P(y|x)\cdot\log Q(y|x)$ (23)
$=\mathop{\mathbb{missing}}{E}\limits_{x\in\mathcal{X}\bowtie\widetilde{\mathcal{I}}}\underbrace{KL(P(y|x)||Q(y||x))}_{\mathrm{point-wise}}+H(\mathcal{Y})+v(\widetilde{\mathcal{I}})+C$ (24)
@@ Appendix C Information quantity relationship in spectral coalitions
Figure 14: Information quantity relationship. This shows the theoretical information quantity relationship between what the characteristic function $v$ measures and the mutual information $\mathbbm{I}(\mathcal{X}\bowtie\widetilde{\mathcal{I}},\mathcal{Y})$. For a given coalition $\widetilde{\mathcal{I}}$, a dataset $\langle\mathcal{X},\mathcal{Y}\rangle$ and a classifier $Q$, the $v$ measures how much information the classifier $Q$ utilizes in decisions. The measured results are then used to compute the marginal contributions of features.
@@ C.2 Normalizing summarized SIDs
=|$=\absolutevalue{\sup||{{\bf\it\beta}}||_{2}\cdot||\bar{\Psi}(v)||_{2}-\frac{||{\bf\it\beta}||_{1}}{M}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{s.t.}\leavevmode\nobreak\ ||\bar{\Psi}(v)||_{1}=1$ (30)
$=\absolutevalue{||{{\bf\it\beta}}||_{2}-\frac{||{\bf\it\beta}||_{1}}{M}}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \mathrm{since}\leavevmode\nobreak\ ||\bar{\Psi}(v)||_{2}^{2}\leq||\bar{\Psi}(v)||_{1}^{2}.$ (31)
@@ C.3 How much samples are sufficient?
(a) resnet18 @CIFAR10
(b) resnet18 @CIFAR100
(c) resnet18 @ImageNet

### v3

@@ 3.2 Spectral coalition filtering
To implement the presences and absences of spectral signals, we define a mask map $\mathbb{T}(\widetilde{\mathcal{I}}):\widetilde{\mathcal{I}}\mapsto\{0,1\}^{M\times N}$ on 2D spectrum, where $\widetilde{\mathcal{I}}$ is a spectral coalition and $M\times N$ denotes 2D image dimensions. The mask map is point-wisely defined as:
$\mathbb{T}(\widetilde{\mathcal{I}})(m,n)=\begin{cases}1,&\text{ifthe frequency point}~{}$(m,n)$~{}\text{is present in coalition}~{}\widetilde{\mathcal{I}},\\ 0,&\text{otherwise}\end{cases}$ (4)
@@
We apply the transfer function $\mathbb{T}(\widetilde{\mathcal{I}})$ on the spectra of images with element-wise product (i.e. Hadamard product (Horn, 1990; Horadam, 2012)). Let $\mathscr{F}$ be the Discrete Fourier transform (DFT) operator and $\mathscr{F}^{-1}$ be the inverse DFT (IDFT) operator (Tan & Jiang, 2018). Readers can further refer to Appendix B.1.
@@
$x\bowtie\widetilde{\mathcal{I}}:=\mathscr{F}^{-1}\left[\underbrace{\mathscr{F}(x)\odot\mathbb{T}(\widetilde{\mathcal{I}})}_{\mathrm{Spectral~{}presence}}+\underbrace{{\bf\it b}\odot(\mathbb{1}-\mathbb{T}(\widetilde{\mathcal{I}}))}_{\mathrm{Spectral~{}absence}}\right]$ (5)
where $\odot$’ denotes Hadamard product (i.e. element-wise product), $\mathbb{1}\in\mathbb{R}^{M\times N}$ denotes an all-ones matrix and ${\bf\it b}\in\mathbb{C}^{M\times N}$ represents the assignments of the absences of spectral players. In the context of attribution analysis, ${\bf\it b}$ is often referred as the baseline. In our implementation, we empirically set ${\bf\it b}={\bf\it 0}$.
@@ 3.3 Characteristic function design
$v(\widetilde{\mathcal{I}})$ $:=\mathop{\mathbb{E}}\limits_{x\thicksim\mathcal{X}}\sum_{y\in\mathcal{Y}}\left\{\underbrace{P(y|x)\cdot\log Q(y|x\bowtie\widetilde{\mathcal{I}})}_{\mathrm{Feature~{}presence}}-\underbrace{P(y|x)\cdot\log Q(y|x\bowtie\emptyset)}_{\mathrm{Absence~{}baseline}}\right\}$
@@
[…]
@@ 3.4 Spectral robustness score (SRS)
The second row in Figure 7 shows the SIDs from various models with randomized weights. We randomize the model weights with Kaiming initialization (He et al., 2015). The measured SIDs exhibit spectral uniformity. This suggests: Un-trained models do not have spectral preferences. We refer to ‘the models with randomized parameters’ as random decisions. Therefore, we assume that the SIDs of random decisions are uniform: $\frac{\mathbb{1}}{M}$.
@@
|$\absolutevalue{{{\bf\it\beta}}^{T}\bar{\Psi}(v)-{{\bf\it\beta}}^{T}\frac{\mathbb{1}}{M}}$ | (10)
where ${{\bf\it\beta}}^{T}\frac{\mathbb{1}}{M}$ is served as a random decision baseline. Let $S(v):v\mapsto[0,1]$ be the normalized result in equation 10. The $S(v)$ is given by:
$S(v):=\frac{\absolutevalue{{{\bf\it\beta}}^{T}\bar{\Psi}(v)-{{\bf\it\beta}}^{T}\frac{\mathbb{1}}{M}}}{\sup\limits_{\bar{\Psi}}\absolutevalue{{{\bf\it\beta}}^{T}\bar{\Psi}(v)-{{\bf\it\beta}}^{T}\frac{\mathbb{1}}{M}}}=\absolutevalue{\frac{{{{\bf\it\bar{\beta}}}}^{T}\bar{\Psi}(v)-\eta}{1-\eta}}$ | (11)
@@ 4.1 Correlation to robustness metrics
[…]
@@ Appendix B Fairness division axioms
Symmetry axiom: Let ~$\widetilde{\mathcal{I}}\in 2^{\mathcal{I}}$ be some spectral player coalition. For $\forall~{}\mathcal{I}_{i},\mathcal{I}_{j}\in\mathcal{I}\land\mathcal{I}_{i},\mathcal{I}_{j}\notin\widetilde{\mathcal{I}}$, the statement $v(\widetilde{\mathcal{I}}\cup\{\mathcal{I}_{i}\})=v(\widetilde{\mathcal{I}}\cup\{\mathcal{I}_{j}\})$ implies $\psi_{i}(\mathcal{I},v)=\psi_{j}(\mathcal{I},v)$. This axiom restates the statement ‘equal treatment of equals’ principle mathematically. This axiom states that the ‘names’ of players should have no effect on the ‘treatments’ by the characteristic function in coalition games (Roth, 1988).
@@
Dummy player axiom: A dummy player (null player) $\mathcal{I}_{*}$ is the player who has no contribution such that: $\psi_{*}(\mathcal{I},v)=0$ and $v(\widetilde{\mathcal{I}}\cup\{\mathcal{I}_{*}\})\equiv v(\widetilde{\mathcal{I}})$ for $\forall~{}\mathcal{I}_{*}\notin\widetilde{\mathcal{I}}\land\mathcal{I}_{*}\subseteq\mathcal{I}$.
@@ B.1 Spectral signal-to-noise ratio (SNR)
$\mathscr{F}(x)(u,v)=\sum_{j=0}^{N-1}\sum_{i=0}^{M-1}x(i,j)e^{-\mathbb{i}2\pi(\frac{u}{M}i+\frac{v}{N}j)}.$ (13)
@@ B.3 Proof for Spectral Coalition Information Identity Theorem
[…]
@@ Appendix C Information quantity relationship in spectral coalitions
Figure 14: Information quantity relationship. This shows the theoretical information quantity relationship between what the characteristic function $v$ measures and the mutual information $\mathbb{I}(\mathcal{X}\bowtie\widetilde{\mathcal{I}},\mathcal{Y})$. For a given coalition $\widetilde{\mathcal{I}}$, a dataset $\langle\mathcal{X},\mathcal{Y}\rangle$ and a classifier $Q$, the $v$ measures how much information the classifier $Q$ utilizes in decisions. The measured results are then used to compute the marginal contributions of features.
@@ C.2 Normalizing summarized SIDs
=|$=\absolutevalue{\sup||{{\bf\it\beta}}||_{2}\cdot||\bar{\Psi}(v)||_{2}-\frac{||{\bf\it\beta}||_{1}}{M}}~{}~{}~{}~{}\mathrm{s.t.}~{}||\bar{\Psi}(v)||_{1}=1$ (30)
$=\absolutevalue{||{{\bf\it\beta}}||_{2}-\frac{||{\bf\it\beta}||_{1}}{M}}~{}~{}~{}~{}\mathrm{since}~{}||\bar{\Psi}(v)||_{2}^{2}\leq||\bar{\Psi}(v)||_{1}^{2}.$ (31)